OpenStudy (anonymous):

Add the complex expressions, and simplify (4+7.9i)+(5+7.3i) can someone help me with this

OpenStudy (anonymous):

You add the "real" numbers with their imaginary; so it is 9 +15.2i

OpenStudy (anonymous):

(4+7.9i)+(5+7.3i) = (4+5) + (7.9+7.3)i = 9 + 15.2i

OpenStudy (anonymous):

Short answer: Simplify the following expression; express in terms of i. square root of -25

OpenStudy (anonymous):

can someone please please help me with this

OpenStudy (anonymous):

remember, even with negatives \[\sqrt{ab} = \sqrt{a}\sqrt{b}\] so let a = 25 , and b = -1 what do you get?

OpenStudy (anonymous):

are you helping with mine

OpenStudy (anonymous):

yes. does this example help? \[\sqrt{-16} = \sqrt{16}\sqrt{-1} = 4 \times \sqrt{-1} = 4i\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

its square root sign with -25 inside

OpenStudy (anonymous):

and says to express in terms of i simplify

OpenStudy (anonymous):

i do not understand how to work this problem can you please help

OpenStudy (anonymous):

how did you get -1 from 16

OpenStudy (anonymous):

dont leave me now